OpenStudy (matt6288):

who wants to help me with slopes?! :D.. formula: y-y1=m(x-X1) solve (11,-9,m=-5

OpenStudy (matt6288):

k so this is what i did. y-(-9)=-5(x-11)

OpenStudy (matt6288):

then y-9-5x+55? +9 +9

OpenStudy (matt6288):

y=-5x+64 so did i do it right?

OpenStudy (anonymous):

This much is right : y - (-9) = -5(x - 11)

OpenStudy (anonymous):

Can you solve left hand side once again??

OpenStudy (anonymous):

y - (-9) = ??

OpenStudy (matt6288):

y-(-9)-5x+55.. i distribute. does that look ok?

OpenStudy (anonymous):

Are you hungry ??? Where = sign have gone?? Eaten ??

OpenStudy (anonymous):

You have distributed right well done.. So you will get now: \[y - (-9) = -5x + 55\]

OpenStudy (matt6288):

ohhh it still stays their :/ no im quite full lol

OpenStudy (matt6288):

ok but after that?

OpenStudy (anonymous):

Tell me : y - (-9) = ??

OpenStudy (matt6288):

y+9=

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

y + 9 = -5x + 55 Subtract 9 both the sides now.

OpenStudy (matt6288):

ohhh ok now i see. y=-5x+46

OpenStudy (anonymous):

Yep..

OpenStudy (matt6288):

can u help me with another lol? (-5/3,7/3),M=6

OpenStudy (matt6288):

y-7/3=6(x-30/3)

OpenStudy (anonymous):

Just do the same steps you will get the answer.. I have to leave now.. I have to go to sleep now its 2:37 am here.. Solve this if you can: \[y - \frac{7}{3} = 6 (x - (-\frac{5}{3}))\]

OpenStudy (matt6288):

y-7/3=6(x-5/3) 1 y-7/3=6x+30/3)?

OpenStudy (anonymous):

Yep go ahead..

OpenStudy (matt6288):

thx! yay i got it lol..

OpenStudy (anonymous):

Okay try it you will get it..

OpenStudy (anonymous):

Good Night.. Take care...

OpenStudy (matt6288):

night